OpenStudy (anonymous):

which binomial is a factor of 4x^2-13x+3? A. 2x-3 B. 2x-1 C. 4x-3 D. 4x-1

OpenStudy (anonymous):

Best to just factorise. However, you could use the factor theorem if you really don't feel like it: If f(a) = 0, than (x-a) is a factor. Note , for example, 2x -1 = 2(x - (1/2))

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you just confused me lol

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

is it d?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

o ok. you didnt confuse me so much

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

Note if you let x = (1/4) \[f(x) = 4 \times \frac{1}{4}^2-13 \times \frac{1}{4} + 3 = 0 \]

OpenStudy (anonymous):

Factors into (4x - 1)(x - 3) = 0